# On Picard’s Theorem Via Nevanlinna Theory

### v5

@@ Abstract.
We consider the Picard’s problem via Nevanlinna theory for complete Kähler manifolds with non-negative Ricci curvature. When the manifold is non-parabolic, which means that it carries a positive global Green function, we give a positive answer to the Picard’s problem. That is, we confirm that every meromorphic function on such a (non-compact) manifold reduces to a constant if it omits three distinct values.
@@
Nevanlinna theory; defect relation; Picard type theorem2010 Mathematics Subject Classification:
32H30; 32H25Contents
@@ 1.1. Motivations and Techniques
It is well-known that the famous Picard’s little theorem asserts that every meromorphic function on the complex Euclidean space reduces to a constant, if it omits three distinct values. All the time, many authors have been trying to generalize this theorem to a more general non-compact complex manifold. Due to S. T. Yau [44], each holomorphic function on a non-compact complete Kähler manifold (or Hermitian manifolds in general) with non-negative Ricci curvature has Liouville property. So, a natural problem is that does Picard’s little theorem still hold on such a manifold? We call it the Picard’s problem. To our knowledge, it is still an unsolved problem. In what follows, we would like to give a precise statement.
Let be a non-compact complete Kähler manifold with non-negative Ricci curvature, see examples for such a manifold in Sha-Yang [34] and Tian-Yau [39, 40]. The well-known Picard’s problem asks that
@@
Remark. Let be the universal covering, where is the unit disc in It is not very hard to confirm that the Picard’s problem can be given a positive answer using Yau’s theorem (see Theorem C below), if any holomorphic mapping lifts to a holomorphic mapping via i.e., Using the criterion for the existence of the liftings, can be lifted to if and only if for some such that Since is simply-connected, then is trivial. This implies that such a lifting exists if and only if is trivial. However, is often not a trivial group, so one cannot lift to in general.
The first result in this direction could be traced back to 1970, S. Kobayashi [20] proved some Picard-type theorem for a holomorphic mappings between complex manifolds from the viewpoint of complex hyperbolicity. As a special consequence, he showed that
Theorem A (Picard’s Theorem, Kobayashi).
Every meromorphic function on on which, a complex Lie group acts transitively, must be a constant if it omits distinct values.
Kobayashi’s result is limited to certain complex manifolds which are acted on transitively by a complex Lie group. In 1975, Goldberg-Ishihara-Petridis [17] presented some Picard-type theorems on harmonic mappings of bounded dilatation on a locally flat Riemannian manifold, refer to N. Petridis [27] also. In particular, they showed that
Theorem B (Picard’s Theorem, Goldberg-Ishihara-Petridis).
Assume that is locally flat. Every holomorphic mapping of bounded dilatation must be a constant mapping.
It does not seem to be a quite satisfactory answer to the Picard’s problem, for some restrictions such as local flatness and bounded dilatation. The first result without any restrictions for is due to S. T. Yau [44], who considered the Liouville’s problem based on the gradient estimation method at the same time, which is viewed as a weaker version of the Picard’s problem. He showed that
Theorem C (Liouville’s Theorem, Yau).
Every holomorphic function on must be a constant if it is bounded.
By S. T. Yau [44], Liouville’s theorem works for all the harmonic functions on (see Cheng-Yau [9] also). The gradient estimation method shows great power in many problems from geometric analysis (see, e.g., [35]), but it seems to be unable to give a solution to the Picard’s problem. Until 2010, A. Atsuji [2] developed a Nevanlinna-type theory based on heat diffusions, and he gave a positive answer to the Picard’s problem for as a meromorphic function on is of a slow growth rate. In the following, we introduce his basic ideas. Let be the Kähler form and Laplace-Beltrami operator of associated to respectively. One can endow with Fubini-Study metric Let be a meromorphic function. Using the Kählerness of the Hilbert-Schmidt norm of differential (with respect to metrics ) can be written as the form
@@
Theorem D (Picard’s Theorem, Atsuji).
@@
where is a geodesic ball centered at a fixed point with radius in Then, must be a constant if it omits distinct values.
Atsuji’s trick [2] (see [11] also) is employing the probabilistic approach (via Brownian motions), who introduced the so-called Nevanlinna-type functions and in which is the time of a Brownian motion on By using Itô’s formula (see, e.g., [18]) and estimates of curvatures, he established a heat diffusion version of Second Main Theorem, which leads to the above Theorem D. It is apparent that is stochastically complete (or is conservative) since has non-negative Ricci curvature by Grigor’yan’s criterion [15]. However, in order to make and meaningful in spirit of Nevanlinna’s settings, the following conditions are necessary:
@@
To do so, has to satisfy certain growth assumptions. That is why a growth condition is needed in Atsuji’s theorem. Unfortunately, the growth condition is so strong that few meromorphic functions can satisfy it.
In the present paper, we shall consider the Picard’s problem by developing Nevanlinna theory (see, e.g., [24, 26, 29]) to complete Kähler manifolds with non-negative Ricci curvature. As we know, Nevanlinna theory studies value distribution of meromorphic mappings between complex spaces, which has been developed for a long time since R. Nevanlinna founded two fundamental theorems for meromorphic functions on the complex plane in 1925. During this period, rich results have been achieved, we may refer the reader to L. V. Ahlfors [4], H. Cartan [6], Carlson-Griffths-King [7, 16], J. Noguchi [23, 24], E. I. Nochka [25], M. Ru [28, 29], B. Shiffman [30], B. Shabat [31], F. Sakai [32, 33], W. Stoll [37, 38], P. Vojta [36], H. Wu [43] and etc., and refer also to [1, 2, 3, 11, 12, 13] and references therein. To our knowledge, almost all the results in Nevanlinna theory are built just on Kähler manifolds which admit a complete Kähler metric of non-positive sectional curvature, while we seem to know little if the domain manifolds are not non-positively curved, because it is hard to estimate the local Green functions for bounded domains under other curvature conditions.
Recently, the author [14] obtained the first result for the value distribution of meromorphic mappings on a complete Kähler manifold with non-negative Ricci curvature by introducing a new method: global Green function method. Based on this method, a positive answer to the Picard’s problem can be given provided that is of maximal volume growth, i.e.,
where denotes the Riemannian volume of geodesic ball centered at a fixed reference point with radius in
The original technique in [14] is to construct a family of relatively compact domains exhausting by using a minimal positive global Green function for Based on the asymptotic estimate of obtained by Colding-Minicozzi [8], we can give an optimal estimate for the local Green function for in terms of integral forms. As a result, the author established the Second Main Theorem of meromorphic mappings on Let us introduce the main contribution and key technique in [14]. Let be the minimal positive global Green function of for Colding-Minicozzi [8] (see [21] also) described the asymptotic behavior of : there exists a constant such that
where denotes the Riemannian distance function of from By means of this asymptotic estimate, the author defined a family of relatively compact domains:
Thus, the Green function of for with a pole at satisfying Dirichlet boundary condition can be written as
@@
holds for all satisfying Based on this estimate, the Calculus Lemma can be derived. When is conformally equivalent to since has non-negative Ricci curvature and maximal volume growth (see [21]). So, by employing the standard arguments, the author established the Second Main Theorem of meromorphic mappings from into a complex projective manifold (see [14], Theorem 3.10), which gives a Picard’s theorem:
Theorem E (Picard’s Theorem, Dong).
Assume that is of maximal volume growth. Every meromorphic function on must be a constant if it omits distinct values.
However, the technique in [14] depends on Colding-Minicozzi’s asymptotic estimate. Unfortunately, there are no asymptotic estimates for if is not of maximal volume growth.
In this paper, we will consider a weaker volume growth condition:
which is equivalent to that carries a positive global Green function. Refer to N. Varopoulos [41, 42] (or see Li-Tam-Wang [19, 21]), there exists uniquely a minimal positive global Green function of for denoted by Our original method is the construction of domain and the application of Li-Yau’s estimates on heat kernels and Green functions.
By Li-Yau’s estimate [22], there exist constants such that
Re-define by
In further, we have
which defines the Green function of for with a pole at satisfying Dirichlet boundary condition. There is a natural relationship between and (see Lemma 4.1): for we have
which plays an important role in the establishment of Nevanlinna theory in this paper.
In the study of Nevanlinna theory, Negative Curvature Method and Logarithmic Derivative Lemma Method are considered as two standard research approaches. In our investigations, Logarithmic Derivative Lemma Method is used. In doing so, it is necessary to establish the so-called Calculus Lemma, which turns out to be a workable road to the Logarithmic Derivative Lemma.
Let us introduce the main techniques: Gradient Estimation (see Theorem 4.2) and Calculus Lemma (see Theorem 5.2) as follows.
Gradient Estimation. For any there exists such that
@@
The above gradient estimate of gives an estimate of the harmonic measure on with respect to : for any there exists such that (see Corollary 4.3)
holds for all
@@
Let be a nonconstant meromorphic function on The Calculus Lemma can lead to Logarithmic Derivative Lemma (see Theorem 5.6): for any there exists a subset of finite Lebesgue measure such that
holds for all outside
@@ 1.2. Main results
Let be a complex projective manifold with over which one can put a Hermitian positive line bundle with Chern form Fix a reduced divisor where denotes the complete linear system of Given a meromorphic mapping we have the Nevanlinna’s functions and see definition in Section 3.2 of this paper. Assume that is non-parabolic, by which we mean that carries a positive global Green function, i.e., it satisfies the volume growth condition
Let be the Ricci form of associated with the metric and be the canonical line bundle over The characteristic function of is defined by
The main result is the following Second Main Theorem.
@@
Let be a non-compact complete Kähler manifold with non-negative Ricci curvature. Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping. If is non-parabolic, then for any there exists a subset of finite Lebesgue measure such that
@@
Again, put
Theorem 1.1 gives a defect relation:
@@
We obtain a Picard’s theorem:
@@
Corollary 1.3 shows that each meromorphic function satisfying on must be a constant, if it omits 2 distinct values.
2. Some Facts from Differential Geometry and Geometric Analysis
@@ 2.1. Volume Comparison Theorem
A space form is defined as a complete Riemannian manifold with constant sectional curvature. Given a simply-connected space form with constant sectional curvature of dimension Let denote the Riemannian volume of a geodesic ball with radius in Besides, let be a complete Riemannian manifold with Ricci curvature of dimension Fix a point Let denote the Riemannian volume of a geodesic ball centered at with radius
The classical volume comparison theorem by Bishop-Gromov (see [5]) can be stated as follows.
@@
Assume that is non-compact. If then we have
holds for all where is the volume of a unit ball in
When Calabi-Yau (see [35]) obtained a lower bound of
@@ 2.2. Estimates of Heat Kernels
Let be a complete Riemannian manifold of dimension with Laplace-Beltrami operator The heat kernel of is the minimal positive fundamental solution of the following heat equation
Fix any reference point Let be the Riemannian distance function of from
@@
If the right-hand integral converges , then defines a minimal positive global Green function of for with a pole at i.e.,
where is the Dirac’s delta function with a pole at It is evident that such a Green function is also unique if (1) is satisfied.
@@
Assume that has non-negative Ricci curvature. If exists, then there exist constants depending only on such that
@@ 3. Nevanlinna’s Functions and First Main Theorem
Let be a non-compact complete Kähler manifold with non-negative Ricci curvature of complex dimension whose Kähler form is defined by
in a local holomorphic coordinate Fix a reference point Let stand for the geodesic ball centered at with radius in Denote by the Riemannian volume of
@@ 3.1. Construction of
Assume that is non-parabolic, which means that satisfies the volume growth condition
Hence, there exists a unique minimal positive global Green function of for i.e.,
where denotes the Laplace-Beltrami operator and is the Dirac function with a pole at Let be the Riemannian distance function of from According to Li-Yau [22], there exist constants such that
Define
By as it is evident to see that is relatively compact for any and the sequence exhausts for According to Sard’s theorem, the boundary of is a submanifold of for almost every Set
which is the positive Green function of for with a pole at satisfying Dirichelet boundary condition, i.e.,
Moreover, we denote by the harmonic measure on with respect to which is defined by
@@ 3.2. Nevanlinna’s Functions
Equip with the Kähler metric Let be the Hilbert-Schmidt norm of differential Since is Kählerian, then we have
Locally, write where is a local holomorphic frame of and is a holomorphic function. By Poincaré-Lelong formula (see, e.g., [7]), it leads to
@@ 3.3. First Main Theorem
To establish the First Main Theorem of we need Jensen-Dynkin formula (see, e.g., [12, 13]) as follows.
@@ 4. Gradient Estimates of Green Functions
Let be a non-compact complete Kähler manifold of complex dimension with non-negative Ricci curvature. Assume that is non-parabolic.
Theorem 4.1.
@@
Let denote the gradient operator on We obtain an estimate of upper bounds of as follows:
@@
On the other hand, we have
Whence, it concludes that

### v6

@@ Abstract.
We study the classical Picard’s problem for complete Kähler manifolds with non-negative Ricci curvature. Based on the global Green function method, we generalize the Nevanlinna theory to complete Kähler manifolds with non-negative Ricci curvature. As a consequence, we give a positive answer to the Picard’s problem in the non-parabolic case, i.e., if the (non-compact) manifold is non-parabolic, then we show that every meromorphic function is a constant if it omits three distinct values.
@@
Nevanlinna theory; meromorphic mappings; Picard’s theorem; defect relation; Kähler manifolds2010 Mathematics Subject Classification:
32H30; 32H25; 32A22Contents
@@ 1.1. Motivations and Techniques
The famous Picard’s theorem asserts that a meromorphic function on the complex Euclidean spaces will reduce to a constant, if it omits three distinct values. Historically, people have been committed to developing this theorem for a long time (see [4, 16, 18, 20, 21, 24, 31]). The study of Picard’s problem on manifolds may have originated from the celebrated work due to S. T. Yau [48] on Liouville’s theorem, which proves the Liouville property of harmonic functions on non-negatively Ricci curved manifolds. Yau’s work has caused a classical problem: does Picard’s theorem hold on complete Kähler manifolds with non-negative Ricci curvature? It is well-known that Nevanlinna theory [30] is a great development of both Liouville’s theorem and Picard’s theorem. So, a further problem is, can Nevanlinna theory extend to complete Kähler manifolds with non-negative Ricci curvature? As far as we know, neither of them has been solved.
Let be a complete non-compact Kähler manifold with non-negative Ricci curvature (see such examples in Sha-Yang [38] and Tian-Yau [43, 44]). The classical Picard’s problem states that
@@
Remark. Let be the universal covering, where is the unit disc in Thanks to the Liouville’s theorem of Yau, we may have an affirmative answer to the Picard’s problem if every holomorphic mapping can lift to a holomorphic mapping by By the mapping lifting theorem, admits a lifting if and only if
for some such that Since is simply-connected, then the fundamental group is trivial. It implies that is necessarily a trivial group if such a lifting exists. However, is not trivial in most cases. Thus, we cannot lift to in general.
An early result in this direction could be dated back to 1970, S. Kobayashi [24] obtained some Picard-type theorems for holomorphic mappings between complex manifolds. In particular, he showed that
Theorem A (Kobayashi, 1970).
If a complex Lie group acts on transitively, then every meromorphic function on must be a constant if it omits distinct values.
Kobayashi’s result is subject to certain complex manifolds which are acted on transitively by a complex Lie group. In 1975, Goldberg-Ishihara-Petridis [21] established some Picard-type theorems on the locally flat manifolds (see N. Petridis [31] also). They showed that
Theorem B (Goldberg-Ishihara-Petridis, 1975).
If is locally flat, then every holomorphic mapping of bounded dilatation must be a constant mapping.
In 2010, A. Atsuji [2] gave an affirmative answer for a class of meromorphic functions of slow growth rate by developing a Nevanlinna-type theory based on heat diffusion, Let us introduce his main result. Let be the Kähler form and Laplace-Beltrami operator of associated to respectively. One can endow with the Fubini-Study metric Let be a meromorphic function. The Kählerness of implies the Hilbert-Schmidt norm of the differential (with respect to ) can be formulated as
@@
Theorem C (Atsuji, 2010).
@@
for any where is the geodesic ball centered at a fixed reference point with radius in Then, must be a constant if it omits distinct values.
Atsuji’s method [2] (see [12] also) is using stochastic calculus by Brownian motion, who introduced the so-called Nevanlinna-type functions and where is the time of the Brownian motion on By computing curvatures and using Itô’s formula (see [22]), he established a heat diffusion version of the Second Main Theorem which leads to the above Theorem C. Note that is stochastically complete (i.e., is conservative), since has non-negative Ricci curvature due to Grigor’yan’s criterion [17]. However, in order to make and meaningful in the sense of Nevanlinna theory, the following conditions are necessary:
@@
To do so, has to satisfy certain growth assumptions. That is why a growth condition is needed in Atsuji’s theorem. Unfortunately, the growth condition seems so strong that there are only a few meromorphic functions satisfy it.
The main purpose of the present paper is to generalize Nevanlinna theory to complete Kähler manifolds with non-negative Ricci curvature. It is known that Nevanlinna theory in several complex variables (see [28, 33]) studies the value distribution of meromorphic mappings between complex spaces, which has been developed for a long time. During this period, fruitful results have been achieved. For example, the reader may refer to L. Ahlfors [5], H. Cartan [7], Carlson-Griffths-King [8, 19], J. Noguchi [27, 28], E. Nochka [29], M. Ru [32, 33], B. Shiffman [34], B. Shabat [35], F. Sakai [36, 37], W. Stoll [41, 42], P. Vojta [40] and H. Wu [47], etc.; and also refer to [1, 2, 3, 12, 13, 14, 15]. To our knowledge, almost all results in Nevanlinna theory are built on complex manifolds with a complete Kähler metric of non-positive sectional curvature, while we seem to know little if the domain manifolds are not non-positively curved, since it is difficult to estimate the local Green function for a general bounded domain, even for a geodesic ball, under other curvature conditions.
Recently, the author [15] gave the first result on the Second Main Theorem in Nevanlinna theory on complete Kähler manifolds with non-negative Ricci curvature by introducing the global Green function method. Consequently, the author answered the Picard’s problem affirmatively provided that is of maximal volume growth:
where denotes the Riemannian volume of a geodesic ball centered at a fixed reference point with radius in
The original technique in [15] is to construct a family of relatively compact exhaustive domains of using the minimal positive global Green function for Based on an asymptotic estimate of obtained by Colding-Minicozzi [9], one can obtain two-sided bounds of the local Green function for in terms of integral forms. As a result, the author established the Second Main Theorem of meromorphic mappings on Let us introduce his main result as follows.
Colding-Minicozzi [9] (see [25] also) described the asymptotic behavior of : there exists a constant such that
where denotes the Riemannian distance function of from By means of the asymptotic estimate, the author defined a family of relatively compact domains as follows
for Hence, the Green function of for with a pole at satisfying Dirichlet boundary condition can be formulated as
@@
holds for all satisfying Based on this estimate, the Calculus Lemma was established. For is conformally equivalent to since has non-negative Ricci curvature and maximal volume growth (see [25]). So, by employing the standard arguments, the author established the Second Main Theorem of meromorphic mappings from into a complex projective manifold (see [15], Theorem 3.10), which gives an immediate result:
Theorem D (Dong, 2023).
If is of maximal volume growth, then every meromorphic function on must be a constant if it omits distinct values.
However, the technique in [15] depends on Colding-Minicozzi’s asymptotic estimate. Unfortunately, there are no such asymptotic estimates for if is not of maximal volume growth.
In this paper, we will generalize the Nevanlinna theory in the case that is a non-parabolic manifold, i.e., there exists a positive global Green function for Note that is non-parabolic if is of maximal volume growth. By N. Varopoulos [45, 46] (see Li-Tam-Wang [23, 25] also), the non-parabolicity implies that there exists a unique minimal positive global Green function of for denoted also by Our original idea is the construction of domain and application of Li-Yau’s estimate on heat kernel and global Green function. Using Li-Yau’s estimate [26], there exist constants such that
holds for all and For we redefine
Then, we have
which defines the Green function of for with a pole at satisfying Dirichlet boundary condition. There is a natural relationship between and : for we have
which plays an important role in the establishment of Nevanlinna theory in this paper (see Lemma 4.1).
Negative Curvature Method and Logarithmic Derivative Lemma Method are very typical approaches in the study of Nevanlinna theory. In this paper, the Logarithmic Derivative Lemma Method is used. So, we need to establish the so-called Calculus Lemma which is a way to the Logarithmic Derivative Lemma. We will introduce two techniques in the paper: Gradient Estimates (see Theorem 4.2) and Calculus Lemma (see Theorem 5.2) as follows.
Gradient Estimates. For any there exists such that
@@
The above gradient estimate yields an estimate of the harmonic measure on with respect to : for any there exists such that
holds for all (see Corollary 4.3).
@@
Let be a nonconstant meromorphic function on The Calculus Lemma can derive the Logarithmic Derivative Lemma: for any there exists a subset of finite Lebesgue measure such that
holds for all outside (see Theorem 5.6).
@@ 1.2. Main results
Let be a complex projective manifold with over which one can put a Hermitian positive line bundle with Chern form We now take a divisor where denotes the complete linear system of Given a meromorphic mapping we have the Nevanlinna’s functions and (see Section 3.2). Let be the Ricci form of associated with the metric and be the canonical line bundle over The characteristic function of is defined by
The main result is the Second Main Theorem as follows.
@@
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping. Then for any there exists a subset of finite Lebesgue measure such that
@@
Set
Since then we obtain a defect relation:
@@
In the non-parabolic case, we answer the Picard’s problem as follows.
@@
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Every meromorphic function on reduces to a constant if it omits distinct values, where denotes the maximal integer not greater than
It implies that every meromorphic function with on reduces to a constant if it omits 2 distinct values.
2. Some Facts from Geometric Analysis
@@ 2.1. Volume Comparison Theorem
A space form is a complete Riemannian manifold with constant sectional curvature. Let be an -dimensional (simply-connected) space form with constant sectional curvature Denote by the Riemannian volume of a geodesic ball with radius in Let be an -dimensional complete Riemannian manifold with Ricci curvature denoted by Denote by the Riemannian volume of the geodesic ball centered at a fixed point with radius in
The well-known volume comparison theorem by Bishop-Gromov (see [6]) states that
@@
We have
holds for all where is the volume of the unit ball in
When Calabi-Yau (see [39]) gave a lower bound of
@@ 2.2. Estimates of Heat Kernels
Let be a complete Riemannian manifold of dimension with Laplace-Beltrami operator Fix a reference point We denote by the Riemannian distance function of from The heat kernel of is the minimal positive fundamental solution to the following heat equation
@@
This infinite integral is convergent if only if is non-parabolic. When is non-parabolic, is the unique minimal positive global Green function of for with a pole at i.e.,
where is the Dirac function with a pole at
When is non-parabolic with non-negative Ricci curvature, Li-Yau [26] (see [39] also) obtained an estimate of as follows.
@@
Assume that is non-parabolic. If has non-negative Ricci curvature, then there exist constants depending only on such that
@@ 3. Nevanlinna’s Functions and First Main Theorem
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature, of complex dimension The Kähler form of associated to the Kähler metric is defined by
@@ 3.1. Construction of
Fix a reference point Let be the Riemannian volume of where is the geodesic ball centered at with radius in Note that the non-parabolicity of implies that
Thus, we have the unique minimal positive global Green function of the half Laplace-Beltrami operator for which can be written as
where is the heat kernel of Let be the Riemannian distance function of from Using Theorem 2.5, there exist constants such that
holds for all For define
By
we see that is a bounded domain containing such that
Furthermore, the family exhausts i.e., for any sequence with we have
Whence, the boundary of can be formulated as
Due to Sard’s theorem, is a submanifold of for almost all
Set
Evidently, is the positive Green function of for with a pole at satisfying Dirichelet boundary condition, i.e.,
Let denote the harmonic measure on with respect to defined by
@@ 3.2. Nevanlinna’s Functions
Equip with the Kähler metric Let be the Hilbert-Schmidt norm of differential with respect to the metrics We have
Locally, write where is a holomorphic local frame of and is a holomorphic function. By Poincaré-Lelong formula (see, e.g., [8]), it leads to
@@ 4. Gradient Estimates of Green Functions
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature, of complex dimension
Lemma 4.1.
@@
Let denote the gradient operator on Then, we have an upper bound of as follows.
@@
Since
then the L’Hospital’rule yields that